# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ Appendix B Achieving the power of a one-sided $t$-test
which is the mid-point of the interval $[v_{-},v_{+}]$, and argued that its numerator is a proxy to a quantity given by,
$-\bm{X}_{j}^{T}(\bm{\hat{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j})\sim\mathcal{N}(-\beta_{j}\bm{X}_{j}^{T}\bm{P}_{-j}\bm{X}_{j},\sigma^{2}\bm{X}_{j}^{T}\bm{P}_{-j}\bm{X}_{j}).$
@@
    Note that for design matrices, $\bm{X}$, of dimension $n\times d$ with i.i.d. drawn Gaussian columns (as is the case with most of the simulations in this paper), it indeed holds that with $d$ getting closer to $n$, the component $\bm{X}_{j}^{T}\bm{P}_{-j}\bm{X}_{j}$ increases in magnitude. Thus, in this case we would expect that the power curves of $\ell$-test and the one-sided $t$-test come closer as $d$ gets closer to $n$ (that is, as we move closer to un-identifiability).
Finally, note that it follows as a direct consequence of the discussions in this section and Section 2.2 that the $\ell$-test gains no power over the $t$-test if $\bm{X}_{j}$ is orthogonal with the rest of the columns (and in particular, for orthogonal designs). In this case, $\hat{m}_{j}=0$, so that we have no estimate of the sign of $\beta_{j}$ and hence, the $\ell$-distribution is symmetric about 0. With smoothing out of the p-value at $1$, we would expect the $\ell$-test and the $t$-test, as well as the respective confidence intervals, to perform similarly.
Appendix C Proofs
C.1 Proof of Lemma 2.1
@@ Appendix C Proofs
which proves (2.1) in the statement of Lemma 2.1. The proof (2.2) follows directly from (Luo et al., 2022, Proposition E.1). $\square$
C.2 Proof of Lemma 2.2
@@
Note that we can write the OLS coefficient as (see (Fithian et al., 2014, Section 4)),
$\hat{\beta}_{j,\mathrm{OLS}}=\frac{\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{y}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}.$ (C.1)
We first start by showing that the OLS estimate, $\hat{\beta}_{j,\mathrm{OLS}}$, is a constant multiple of the statistic, $u_{1}$, where the constant is a deterministic function of the sufficient statistic, $\bm{S}^{(j)}$. Based on Lemma 2.1, we have the decomposition,
@@
$\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{y}&=\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{P}_{-j}\bm{y}+\hat{\sigma}_{j}\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{V}\bm{u}\\ &=\hat{\sigma}_{j}\bm{X}_{j}^{T}\bm{V}\bm{u}.\textrm{ [since the columns of }\bm{V}\textrm{ are orthogonal to the columns of }\bm{X}_{-j}]\end{aligned}$ (C.2)
@@
$\hat{\beta}_{j,\mathrm{OLS}}=\frac{\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\hat{\bm{e}}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}.=\frac{\hat{\sigma}_{j}u_{1}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|}$ (C.4)
@@
$&\hat{\bm{e}}_{\parallel}=\frac{(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\hat{\bm{e}}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}=(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\hat{\beta}_{j,\mathrm{OLS}}\textrm{, and }\\ &\hat{\bm{e}}_{\perp}=\left(\bm{I}-\frac{(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}\right)\hat{\bm{e}}_{j}.\end{aligned}$ (C.5)
@@
C.3 Proof of Theorem 2.1
We will first prove that $f_{\bm{S}^{(j)}}$ is continuous and strictly increasing in $\{u:f_{\bm{S}^{(j)}}(u)\neq 0\}$ (in C.3.1), followed by a proof of the characterization of its inverse (in C.3.2).
C.3.1 Proof of continuity and increasing properties of $f_{\mathbf{S}_{j}}$
@@
and as we did in the proof in Section C.2, we again decompose,
@@
Next using Equation (C.1) and (C.4),
@@
Also from Equation (C.5),
@@
$&\|\bm{y}-\bm{X}\bm{\beta}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}+\hat{\bm{e}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{X}_{j}\beta_{j}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}+\hat{\bm{e}}_{j}-(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\beta_{j}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}+\hat{\bm{e}}_{\perp}+\hat{\bm{e}}_{\parallel}-(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\beta_{j}\|^{2}\\ &+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\textrm{ [}\hat{\bm{e}}_{\perp},\hat{\bm{e}}_{\parallel}\textrm{ are defined above]}\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}\|^{2}+\|\hat{\bm{e}}_{\perp}\|^{2}+\|\hat{\bm{e}}_{\parallel}-(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\beta_{j}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}\|^{2}+\|\hat{\bm{e}}_{\perp}\|^{2}+\left(\frac{u_{1}\hat{\sigma}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|}-\beta_{j}\right)^{2}\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}\\ &+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\tilde{f}(\bm{\beta};\bm{y},\bm{X})+\left(\frac{u_{1}\hat{\sigma}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|}-\beta_{j}\right)^{2}\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2},\end{aligned}$ (C.6)
@@
$\hat{\bm{\beta}}(a):=\underset{\bm{\beta}}{\arg\min}\left(\underbrace{\tilde{f}+\left(a-\beta_{j}\right)^{2}\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}_{=:U(a,\bm{\beta};\bm{y},\bm{X})}\right).$ (C.7)
Thus, $\hat{\beta}_{j}\left(\frac{\hat{\beta}_{j,\mathrm{OLS}}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}\right)$ is just the LASSO estimate of $\beta_{j}$, $\hat{\beta}^{\lambda}_{j}$ and hence, it suffices to show that $\hat{\beta}^{\lambda}_{j}(a)$ is a non-decreasing function of $a$. As is the case with the function, $\tilde{f}$, we will also be primarily be interested in the behavior of $U$ as a function of the arguments, $(a,\bm{\beta})$. We start with listing some properties of the function, $U$. First, note that $U$ is a continuous, convex function in its arguments and is strictly convex in $a$ for any fixed value of $\bm{\beta}$. This implies that $\hat{\bm{\beta}}(a)$ is continuous in $a$. To see this, first note that as $\|\bm{\beta}\|\rightarrow\infty$, the function, $U$ diverges, so that for the minimization problem to obtain $\hat{\bm{\beta}}(a)$, we can constraint $\bm{\beta}$ within some compact set, $C(a)\subset\mathbb{R}^{d}$. One can then apply Berge’s Maximum Theorem (Berge, 1963), to conclude that $\hat{\bm{\beta}}(a)$ is continuous in $a$.
@@
Note that because $\hat{\bm{\beta}}(a)$ maximizes $U$, and $\forall i\notin\mathcal{A}(a)\cup\{j\}$, $\hat{\bm{\beta}}_{j}(a)=0$, it holds that $\hat{\bm{\beta}}_{\mathcal{A}(a)\cup\{j\}}(a)$ is a minimizer of $U(a,\bm{\gamma};\bm{y},\bm{X}_{\mathcal{A}(a)\cup\{j\}})$, where now $\bm{\gamma}$ is a vector of length $|\mathcal{A}(a)\cup\{j\}|$. For a proof, see Lemma 1 of (Liu et al., 2021+, Section 3.2). Because all the entries of $\hat{\bm{\beta}}_{\mathcal{A}(a)\cup\{j\}}$ are non-zero, $U(a,\bm{\gamma};\bm{y},\bm{X}_{\mathcal{A}(a)\cup\{j\}})$ is differentiable in $\bm{\gamma}$ at $\hat{\bm{\beta}}_{\mathcal{A}(a)\cup\{j\}}$, and because the latter is a minimizer, an appeal to the first-order stationary conditions yield that for any $i\in\mathcal{A}(a)$,
@@
Note that from the continuity of $\hat{\bm{\beta}}(a)$ in the neighbourhood of $a$ where $\mathcal{A}(a)$ does not change, the sign of the active variables also remain constant so that $\mathrm{sign}(\hat{\beta}_{i}(a))$ is a constant in that neighborhood, $\forall i\in\mathcal{A}(a)\cup\{j\}$. Thus, differentiating the above equation both the sides with respect to $a$ yields,

### v2

@@ Appendix B Achieving the power of a one-sided $t$-test
which is the mid-point of the interval $v_{-},v_{+}]$, and argued that its numerator is a proxy to a quantity given by,
$\bm{X}_{j}^{T}(\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j})\sim\mathcal{N}(-\beta_{j}\bm{X}_{j}^{T}\bm{P}_{-j}\bm{X}_{j},\sigma^{2}\bm{X}_{j}^{T}\bm{P}_{-j}\bm{X}_{j}).$
@@
    Note that for design matrices, $\bm{X}$, of dimension $n\times d$ with i.i.d. drawn Gaussian columns (as is the case with most of the simulations in this paper), it indeed holds that with $d$ getting closer to $n$, the component $\bm{X}_{j}^{T}\bm{P}_{-j}\bm{X}_{j}$ increases in magnitude. Thus, in this case we would expect that the power curves of $\ell$-test and the one-sided $t$-test come closer as $d$ gets closer to $n$ (that is, as we move closer to unidentifiability).
Finally, note that it follows as a direct consequence of the discussions in this section and Section 2.2 that the $\ell$-test gains no power over the $t$-test if $\bm{X}_{j}$ is orthogonal with the rest of the columns (and in particular, for orthogonal designs). In this case, $\hat{m}_{j}=0$, so that we have no estimate of the sign of $\beta_{j}$ and hence, the $\ell$-distribution is symmetric about 0. With smoothing out of the $p$-value at 11, we would expect the $\ell$-test and the $t$-test, as well as the respective confidence intervals, to perform similarly.
Appendix C $\ell$-test when $\sigma$ is known
The $\ell$-test can easily be extended to the case with a known error variance $\sigma$, in particular, with the following substitutions over the usual, unknown-$\sigma$-case results: For a generic $(\bm{y},\bm{X}_{n\times d})$—the sufficient statistic under $H_{j}$ now just shrinks to its first element and, re-using the old notation, is now given by $\bm{S}^{(j)}=\bm{X}_{-j}^{T}\bm{y}$, while the known-$\sigma$ analogue of Equation (2.2) is
$\bm{y}=\bm{P}_{-j}\bm{y}+\sigma\bm{V}\bm{z},$ (C.1)
for some unique (nd+1)(n-d+1)-dimensional vector $\bm{z}$ with distribution $\mathcal{N}(\bm{0},\bm{I}_{n-d+1})$ under $H_{j}$. A known-$\sigma$ analogue of Theorem 1 still holds true with $\hat{\sigma}_{j}$ replaced by the known $\sigma$ and $u_{1}$ replaced by $z_{1}$, and in particular the known-$\sigma$ analogue of Equation (2.5) equates the event $\{\hat{\beta}_{j}^{\lambda}\leq b\}$ with $\left\{z_{1}\leq\Lambda_{j}(b,\mathrm{sign}(b))\right\}$, where $\Lambda_{j}(b,\mathrm{sign}(b))$ is easily seen to be a function only of ${\bm{S}}^{(j)}$ as required for a straightforward evaluation of the probability of this event via the Gaussian CDF, since the known-$\sigma$ expression of $\Lambda_{j}(b,\mathrm{sign}(b))$ (defined by the aforementioned known-$\sigma$ analogue of Theorem 1) now has $\sigma$, instead of $\hat{\sigma}_{j}$ in the denominator. The known-$\sigma$ analogues of Sections 2.3 and 2.4 similarly replace $\bm{u}$ and $\tilde{\bm{u}}$ with $\bm{z}$ and $\tilde{\bm{z}}$, respectively, along with replacing their corresponding distributions (in particular, Unif($\mathbb{S}^{n-d}$) is replaced with $\mathcal{N}(\bm{0},\bm{I}_{n-d+1})$).
In Section 3, we discuss a re-centered test, which, in the unknown-$\sigma$-case, would be based on the test statistic $|u_{1}-\hat{m}_{j}|$, but in the known-$\sigma$ case (which is the case considered in Section 3 and Theorem 3 therein) the re-centered test replaces $u_{1}$ with $z_{1}$ and replaces $\hat{\sigma}_{j}$ in the expression for $\hat{m}_{j}$ with the known $\sigma$. We will denote this known-$\sigma$ analogue of $\hat{m}_{j}$ by $\hat{m}_{j}^{\prime}$.
Appendix D Proofs
D.1 Proof of Lemma 1
@@
which proves (2.1) in the statement of Lemma 1. The proof (2.2) follows directly from luo2022. $\square$
D.2 Proof of Lemma 2
@@
Note that we can write the OLS coefficient as (see fithianthesis),
$\hat{\beta}_{j,\mathrm{OLS}}=\frac{\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{y}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}.$ (D.1)
We first start by showing that the OLS estimate, $\hat{\beta}_{j,\mathrm{OLS}}$, is a constant multiple of the statistic, $u_{1}$, where the constant is a deterministic function of the sufficient statistic, $\bm{S}^{(j)}$. Based on Lemma 1, we have the decomposition,
@@
[…]
@@
$\hat{\beta}_{j,\mathrm{OLS}}=\frac{\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\hat{\bm{e}}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}.=\frac{\hat{\sigma}_{j}u_{1}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|}$ (D.4)
@@
$\begin{aligned} &\hat{\bm{e}}_{\parallel}=\frac{(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\hat{\bm{e}}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}=(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\hat{\beta}_{j,\mathrm{OLS}}\textrm{, and }\\ &\hat{\bm{e}}_{\perp}=\left(\bm{I}-\frac{(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}\right)\hat{\bm{e}}_{j}.\end{aligned}$ (D.5)
@@
D.3 Proof of Theorem 1
We first start with an auxiliary lemma that states conditions to establish the continuity (in its remaining arguments) of the minimizer of a marginal of a function. In order to do so, we first recall the concept of upper hemicontinuity of a set-valued function.
Definition D.1 (Upper hemicontinutiy).
Consider topological spaces $X$ and $Y$ and consider a set-valued function $\phi:X\mapsto 2^{Y}$. We call $\phi$ upper hemicontinuous at a point $x\in X$, if for any open set $U\subseteq Y$ such that $\phi(x)\subseteq U$, there exists a neighborhood $V$ of $x$ such that $w\in V\implies\phi(w)\subseteq U$.
It is easy to see that if the set-valued function $\phi$ is actually a function (that is, only singleton sets belong to its range), then hemicontinuity of $\phi$ at $x$ actually implies the continuity of $\phi$ at $x$ as a function, as the definition implies that pre-image under $\phi$ of every open set is open. Now we state the following lemma that will be used multiple times in this, and some of the following proofs.
Lemma 3.
Consider a real-valued function $\phi:\mathcal{X}\times\mathcal{Y}\mapsto\mathbb{R}_{\geq 0}$ (with $\mathcal{X},\mathcal{Y}$ subsets of Euclidean spaces) and $\bm{y}^{*}\in\mathcal{Y}$ satisfying the following properties:
  1. 1.
    $(\bm{x},\bm{y})\mapsto\phi(\bm{x},\bm{y})$ is continuous.
  2. 2.
    For all $d>0$, there exists $M:=M(d)>0$ such that for $\epsilon=\underset{\bm{y}\in\bar{B}(\bm{y}^{*},d)}{\sup}\phi(\bm{0},\bm{y})$ (this supremum exists as $\bm{y}\mapsto\phi(\bm{0},\bm{y})$ is a continuous transformation and $\bar{B}(\bm{y}^{*},d)=\{\bm{y}:\|\bm{y}-\bm{y}^{*}\|\leq d\}$ is a closed ball), we have that $\|\bm{x}\|>M\implies\underset{\bm{y}\in\bar{B}(\bm{y}^{*},d)}{\inf}\phi(\bm{x},\bm{y})>\epsilon$.
Then the map $\bm{y}\mapsto\underset{\bm{x}}{\min}~\phi(\bm{x},\bm{y})$ is continuous at $\bm{y}^{*}$ and the map $\bm{y}\mapsto\underset{\bm{x}}{\arg\min}~\phi(\bm{x},\bm{y})$ is upper hemicontinuous (where $\arg\min$ denotes the set of possible minimizers) at $\bm{y}^{*}$. Hence, if the minimizer is unique for values of $\bm{y}$ in a neighborhood of $\bm{y}^{*}$, then the map is continuous at $\bm{y}^{*}$.
Proof.
Fix any $\bm{y}^{*}\in\mathcal{Y}$ and let $G\subset\mathcal{Y}$ be a compact set containing $\bm{y}^{*}$. Due to the Hiene–Borel theorem (rudin1976principles, Theorem 2.41), $G$ is closed and bounded and hence, there is some closed ball $\bar{B}(\bm{y}^{*},d)$ of radius $d>0$ centerd at $\bm{y}^{*}$ such that $G\subset\bar{B}(\bm{y}^{*},d)$. Note that due to Property 2 above we have that with $M=M(d)$, $\|\bm{x}\|>M\implies\phi(\bm{x},\bm{y})>\underset{\bm{y}\in\bar{B}(\bm{y}^{*},d)}{\sup}\phi(\bm{0},\bm{y})$ for all $\bm{y}\in\bar{B}(\bm{y}^{*},d)$. This implies that for any $\bm{y}\in\bar{B}(\bm{y}^{*},d)$, $\phi(\bm{0},\bm{y})<\phi(\bm{x},\bm{y})$ for any $\|\bm{x}\|>M$. This implies that for any $\bm{y}\in\bar{B}(\bm{y}^{*},d)$,
$\underset{\bm{x}}{\arg\min}~\phi(\bm{x},\bm{y})=\underset{\bm{x}:\|\bm{x}\|\leq M}{\arg\min}~\phi(\bm{x},\bm{y})$
Now define the set-valued function:
$C(\bm{y})=\{\bm{x}:\|\bm{x}\|\leq M\},$
and note that $C(\bm{y})$ is compact and $\bm{y}\mapsto C(\bm{y})$ is a constant function. We can then apply Berge’s Maximum Theorem (berge) to conclude that the map $\bm{y}\mapsto\underset{\bm{x}\in C(\bm{y})}{\arg\min}~\phi(\bm{x},\bm{y})$ is upper hemicontinuous at $\bm{y}^{*}$, and since that map is equal to the map $\underset{\bm{x}}{\arg\min}~\phi(\bm{x},\bm{y})$ in a neighborhood of $\bm{y}^{*}$, the latter map is also upper hemicontinuous at $\bm{y}^{*}$. Similarly, Berge’s Maximum Theorem implies the map $\bm{y}\mapsto\underset{\bm{x}\in C(\bm{y})}{\min}~\phi(\bm{x},\bm{y})$ is continuous at $\bm{y}^{*}$, and since that map is equal to the map $\underset{\bm{x}}{\min}~\phi(\bm{x},\bm{y})$ in a neighborhood of $\bm{y}^{*}$, the latter map is also continuous at $\bm{y}^{*}$. $\square$
Next, we first prove that $f_{\bm{S}^{(j)}}$ is continuous and strictly increasing in $\{u:f_{\bm{S}^{(j)}}(u)\neq 0\}$ (in D.3.1), followed by a proof of the characterization of its inverse (in D.3.2).
D.3.1 Proof of continuity and increasing properties of $f_{\mathbf{S}_{j}}$
@@
and as we did in the proof in Section D.2, we again decompose,
@@
Next using Equation (D.1) and (D.4),
@@
Also from Equation (D.5),
@@
$\begin{aligned} &\|\bm{y}-\bm{X}\bm{\beta}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}+\hat{\bm{e}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{X}_{j}\beta_{j}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}+\hat{\bm{e}}_{j}-(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\beta_{j}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}+\hat{\bm{e}}_{\perp}+\hat{\bm{e}}_{\parallel}-(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\beta_{j}\|^{2}\\ &+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\textrm{ [}\hat{\bm{e}}_{\perp},\hat{\bm{e}}_{\parallel}\textrm{ are defined above]}\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}\|^{2}+\|\hat{\bm{e}}_{\perp}\|^{2}+\|\hat{\bm{e}}_{\parallel}-(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\beta_{j}\|^{2}+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\|\hat{\bm{y}}_{j}-\bm{X}_{-j}\bm{\beta}_{-j}-\bm{P}_{-j}\bm{X}_{j}\beta_{j}\|^{2}+\|\hat{\bm{e}}_{\perp}\|^{2}+\left(\frac{u_{1}\hat{\sigma}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|}-\beta_{j}\right)^{2}\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}\\ &+2n\lambda|\beta_{j}|+2n\lambda\sum_{i\neq j}|\beta_{i}|\\ &=\tilde{f}(\bm{\beta};\bm{y},\bm{X})+\left(\frac{u_{1}\hat{\sigma}_{j}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|}-\beta_{j}\right)^{2}\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2},\end{aligned}$ (D.6)
@@
$\hat{\bm{\beta}}(a):=\underset{\bm{\beta}}{\arg\min}\left(\underbrace{\tilde{f}+\left(a-\beta_{j}\right)^{2}\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}_{=:U(a,\bm{\beta};\bm{y},\bm{X})}\right).$ (D.7)
Thus, $\hat{\beta}_{j}\left(\frac{\hat{\beta}_{j,\mathrm{OLS}}}{\|(\bm{I}-\bm{P}_{-j})\bm{X}_{j}\|^{2}}\right)$ is just the LASSO estimate of $\beta_{j}$, $\hat{\beta}^{\lambda}_{j}$ and hence, it suffices to show that $\hat{\beta}^{\lambda}_{j}(a)$ is a non-decreasing function of $a$. As is the case with the function, $\tilde{f}$, we will also be primarily be interested in the behavior of $U$ as a function of the arguments, $(a,\bm{\beta})$. We start with listing some properties of the function, $U$.
First, note that $U$ is a continuous, convex function in its arguments and is strictly convex in $a$ for any fixed value of $\bm{\beta}$. Next, we will show that $\hat{\bm{\beta}}(a)$ is continuous in $a$ by invoking Lemma 3. In order to do so, observe that $(a,\bm{\beta})\mapsto U(a,\bm{\beta})$ is a continuous function satisfying item 1 of the requirements. Next, observe that $U(a,\bm{\beta})\geq\tilde{f}(\bm{\beta})\geq 2n\lambda\|\bm{\beta}\|_{1}$ for any $a$ and $\bm{\beta}$, and note that $2n\lambda\|\bm{\beta}\|_{1}$ does not depend on $a$. Thus, noting that $\lambda>0$, for any $d$ and $a^{*}$, we have that $\inf_{|a-a^{*}|\leq d}U(a,\bm{\beta})\geq 2n\lambda\|\bm{\beta}\|_{1}\to\infty$, as $\|\bm{\beta}\|\to\infty$. This satisfies item 2, and we can now invoke Lemma 3 to conclude the continuity of $a\mapsto\hat{\bm{\beta}}(a)$.
@@
Note that because $\hat{\bm{\beta}}(a)$ minimizes $U$, and $\forall i\notin\mathcal{A}(a)\cup\{j\}$, $\hat{\bm{\beta}}_{j}(a)=0$, it holds that $\hat{\bm{\beta}}_{\mathcal{A}(a)\cup\{j\}}(a)$ is a minimizer of $U(a,\bm{\gamma};\bm{y},\bm{X}_{\mathcal{A}(a)\cup\{j\}})$, where now $\bm{\gamma}$ is a vector of length $|\mathcal{A}(a)\cup\{j\}|$. For a proof, see Lemma 1 of dcrt. Because all the entries of $\hat{\bm{\beta}}_{\mathcal{A}(a)\cup\{j\}}$ are non-zero, $U(a,\bm{\gamma};\bm{y},\bm{X}_{\mathcal{A}(a)\cup\{j\}})$ is differentiable in $\bm{\gamma}$ at $\hat{\bm{\beta}}_{\mathcal{A}(a)\cup\{j\}}$, and because the latter is a minimizer, an appeal to the first-order stationary conditions yield that for any $i\in\mathcal{A}(a)$,
@@
Note that from the continuity of $\hat{\bm{\beta}}(a)$ in the neighborhood of $a$ where $\mathcal{A}(a)$ does not change, the sign of the active variables also remain constant so that $\mathrm{sign}(\hat{\beta}_{i}(a))$ is a constant in that neighborhood, $\forall i\in\mathcal{A}(a)\cup\{j\}$. Thus, differentiating the above equation both the sides with respect to $a$ yields,